A snooker ball X of mass 0.3kg moving with a velocity 5m/s hit a stationary ball of mass 0.4kg.Y moves off with a velocity of 2m/s at 30degree to the initial direction of X.Find the velocity v of X and its direction after hitting Y.

1
Expert's answer
2022-01-27T09:12:18-0500

Apply conservation of momentum. The momentum is conserved along x-axis (along the initial direction of motion of X) and along y-axis (perpendicular to the stated x-axis):


"Ox: m_xv_x=m_yu_y\cos\theta+m_xu_x\cos\phi,\\
Oy: 0=m_yu_y\sin\theta+m_xu_x\sin\phi.\\\space\\
m_xu_x\cos\phi=m_xv_x-m_yu_y\cos\theta,\\
m_xu_x\sin\phi=-m_yu_y\sin\theta.\\\space\\
\tan\phi=\dfrac{m_yu_y\sin\theta}{m_yu_y\cos\theta-m_xv_x}=-0.5,\\\space\\
\phi=-26.6°,\\
u_x=2.98\text{ m/s}."